# Adaptive Extraction Network for Multivariate Long Sequence Time-Series Forecasting

### v1

@@ 4.4 Multivariate results and analysis
ETTm1 96 0.142 0.263 0.264 0.330 0.320 0.357 0.334 0.368 0.319 0.358 0.324 0.361 0.316 0.362 0.338 0.375 0.428 0.444 0.345 0.372 0.375 0.398 0.379 0.419 0.386 0.398 0.806 0.657 0.450 0.477
192 0.162 0.282 0.316 0.368 0.361 0.381 0.377 0.391 0.369 0.384 0.362 0.383 0.363 0.390 0.374 0.387 0.445 0.468 0.380 0.389 0.408 0.410 0.426 0.441 0.459 0.444 0.923 0.711 0.469 0.481
336 0.174 0.295 0.315 0.378 0.390 0.404 0.426 0.420 0.402 0.406 0.390 0.402 0.408 0.426 0.410 0.411 0.533 0.519 0.413 0.413 0.435 0.428 0.445 0.459 0.495 0.464 1.062 0.778 0.586 0.574
720 0.184 0.304 0.382 0.425 0.454 0.441 0.491 0.459 0.464 0.442 0.461 0.438 0.481 0.476 0.478 0.450 0.728 0.655 0.474 0.453 0.499 0.462 0.543 0.490 0.585 0.516 1.208 0.853 0.635 0.597
Avg 0.166 0.286 0.319 0.375 0.381 0.395 0.407 0.410 0.389 0.398 0.384 0.396 0.392 0.414 0.400 0.406 0.534 0.522 0.403 0.407 0.429 0.425 0.448 0.452 0.481 0.456 1.000 0.750 0.535 0.532
ETTm2 96 0.111 0.219 0.183 0.265 0.175 0.258 0.180 0.264 0.175 0.258 0.177 0.260 0.179 0.275 0.187 0.267 0.197 0.321 0.193 0.292 0.189 0.280 0.203 0.287 0.192 0.274 1.046 0.810 0.243 0.342
192 0.129 0.236 0.248 0.305 0.237 0.299 0.250 0.309 0.240 0.230 0.248 0.306 0.307 0.376 0.249 0.309 0.326 0.375 0.284 0.362 0.253 0.319 0.269 0.328 0.280 0.339 2.049 1.108 0.392 0.448
336 0.140 0.247 0.311 0.345 0.298 0.340 0.311 0.348 0.301 0.338 0.304 0.342 0.325 0.388 0.321 0.351 0.372 0.421 0.369 0.427 0.314 0.357 0.325 0.366 0.334 0.361 2.298 1.199 0.932 0.724
720 0.152 0.255 0.101 0.396 0.391 0.396 0.412 0.407 0.404 0.398 0.403 0.397 0.502 0.490 0.408 0.403 0.410 0.448 0.554 0.522 0.414 0.413 0.421 0.415 0.417 0.413 2.556 1.317 1.372 0.879
Avg 0.133 0.239 0.211 0.328 0.275 0.323 0.288 0.332 0.280 0.306 0.283 0.326 0.328 0.382 0.291 0.333 0.326 0.391 0.350 0.401 0.293 0.342 0.305 0.349 0.306 0.347 1.987 1.109 0.735 0.598
ETTh1 96 0.203 0.326 0.338 0.368 0.375 0.400 0.386 0.405 0.381 0.391 0.394 0.408 0.398 0.427 0.384 0.402 0.429 0.440 0.386 0.400 0.494 0.479 0.376 0.419 0.513 0.491 0.602 0.546 0.548 0.528
192 0.228 0.347 0.377 0.385 0.429 0.421 0.441 0.436 0.433 0.420 0.446 0.438 0.430 0.453 0.436 0.429 0.494 0.482 0.437 0.432 0.538 0.504 0.420 0.448 0.534 0.504 0.663 0.584 0.542 0.526
336 0.238 0.355 0.400 0.404 0.484 0.458 0.487 0.458 0.472 0.441 0.485 0.455 0.460 0.460 0.491 0.469 0.706 0.625 0.481 0.459 0.574 0.521 0.459 0.465 0.588 0.535 0.729 0.622 1.298 0.942
720 0.245 0.363 0.465 0.439 0.498 0.482 0.503 0.491 0.485 0.471 0.495 0.474 0.491 0.509 0.521 0.500 0.750 0.689 0.519 0.516 0.562 0.535 0.506 0.507 0.643 0.616 0.846 0.685 0.721 0.659
Avg 0.229 0.348 0.395 0.399 0.447 0.440 0.454 0.448 0.443 0.431 0.455 0.444 0.445 0.462 0.458 0.450 0.595 0.559 0.456 0.452 0.542 0.510 0.440 0.460 0.570 0.537 0.710 0.609 0.777 0.664
ETTh2 96 0.147 0.259 0.385 0.396 0.289 0.341 0.297 0.349 0.289 0.338 0.302 0.348 0.332 0.377 0.340 0.374 0.632 0.547 0.333 0.387 0.340 0.391 0.358 0.397 0.476 0.458 0.380 0.460 1.616 1.036
192 0.168 0.279 0.438 0.425 0.372 0.392 0.380 0.400 0.375 0.391 0.388 0.400 0.422 0.441 0.402 0.414 0.876 0.663 0.477 0.476 0.430 0.439 0.429 0.439 0.512 0.493 0.574 0.561 2.083 1.197
336 0.192 0.295 0.451 0.437 0.386 0.414 0.428 0.432 0.425 0.435 0.426 0.433 0.447 0.474 0.452 0.452 0.924 0.702 0.594 0.541 0.485 0.479 0.496 0.487 0.552 0.551 0.973 0.932 2.970 1.439
720 0.223 0.310 0.470 0.461 0.412 0.434 0.427 0.445 0.435 0.449 0.431 0.446 0.442 0.467 0.462 0.468 1.390 0.863 0.831 0.657 0.500 0.497 0.463 0.474 0.562 0.560 2.661 1.257 2.576 1.363
Avg 0.183 0.286 0.436 0.430 0.364 0.395 0.383 0.407 0.381 0.403 0.387 0.407 0.411 0.440 0.414 0.427 0.956 0.694 0.559 0.515 0.439 0.452 0.437 0.449 0.526 0.516 1.147 0.803 2.311 1.259
Electricity 96 0.131 0.233 0.179 0.263 0.153 0.247 0.148 0.240 0.146 0.244 0.195 0.285 0.164 0.269 0.168 0.272 0.254 0.347 0.197 0.282 0.187 0.304 0.193 0.308 0.169 0.273 0.312 0.395 0.300 0.392
192 0.135 0.239 0.185 0.269 0.166 0.256 0.162 0.253 0.163 0.259 0.199 0.289 0.177 0.285 0.184 0.289 0.261 0.353 0.196 0.285 0.199 0.315 0.201 0.315 0.182 0.286 0.347 0.425 0.297 0.390
336 0.138 0.245 0.201 0.285 0.185 0.277 0.178 0.269 0.180 0.279 0.215 0.305 0.193 0.304 0.198 0.300 0.273 0.364 0.209 0.301 0.212 0.329 0.214 0.329 0.200 0.304 0.351 0.439 0.317 0.403
720 0.142 0.250 0.242 0.318 0.225 0.310 0.225 0.317 0.216 0.307 0.256 0.337 0.212 0.321 0.220 0.320 0.303 0.388 0.245 0.333 0.233 0.345 0.246 0.355 0.222 0.321 0.382 0.448 0.338 0.417
Avg 0.137 0.242 0.202 0.284 0.182 0.272 0.178 0.270 0.176 0.272 0.216 0.304 0.187 0.295 0.193 0.295 0.273 0.363 0.212 0.300 0.208 0.323 0.214 0.327 0.193 0.296 0.348 0.427 0.313 0.401
Traffic 96 0.228 0.286 0.376 0.265 0.462 0.285 0.395 0.268 0.483 0.321 0.544 0.359 0.519 0.309 0.593 0.321 0.558 0.320 0.650 0.396 0.607 0.392 0.587 0.366 0.612 0.338 0.999 0.596 0.798 0.436
192 0.239 0.292 0.392 0.271 0.473 0.296 0.417 0.276 0.490 0.321 0.540 0.354 0.537 0.315 0.617 0.336 0.572 0.331 0.598 0.370 0.621 0.399 0.604 0.373 0.613 0.340 1.008 0.597 0.849 0.481
336 0.243 0.300 0.405 0.277 0.498 0.296 0.433 0.283 0.506 0.330 0.551 0.358 0.534 0.313 0.629 0.336 0.587 0.342 0.605 0.373 0.622 0.396 0.621 0.383 0.618 0.328 1.034 0.602 0.828 0.476
720 0.256 0.307 0.436 0.294 0.506 0.313 0.467 0.302 0.538 0.341 0.586 0.375 0.577 0.325 0.640 0.350 0.652 0.359 0.645 0.394 0.632 0.396 0.626 0.382 0.653 0.355 - - 0.854 0.489
Avg 0.242 0.296 0.402 0.277 0.484 0.297 0.428 0.282 0.504 0.328 0.555 0.362 0.542 0.316 0.620 0.336 0.592 0.338 0.625 0.383 0.621 0.396 0.610 0.376 0.624 0.340 1.014 0.598 0.832 0.471
Exchange 96 0.097 0.083 0.083 0.196 0.100 0.222 0.086 0.206 0.082 0.199 0.082 0.201 0.102 0.235 0.107 0.234 0.329 0.440 0.088 0.218 0.085 0.204 0.148 0.237 0.111 0.237 0.516 0.529 0.395 0.474
192 0.103 0.170 0.170 0.291 0.212 0.326 0.177 0.299 0.176 0.297 0.187 0.307 0.172 0.316 0.226 0.344 0.544 0.586 0.176 0.315 0.182 0.303 0.271 0.335 0.219 0.335 0.97 0.764 0.776 0.698
336 0.117 0.316 0.316 0.404 0.379 0.442 0.331 0.417 0.337 0.418 0.345 0.427 0.272 0.407 0.367 0.448 1.017 0.786 0.313 0.427 0.348 0.428 0.460 0.476 0.421 0.476 1.766 1.039 1.029 0.797
720 0.158 0.817 0.817 0.679 0.904 0.715 0.847 0.691 0.909 0.725 0.887 0.708 0.714 0.658 0.964 0.746 1.239 0.912 0.839 0.695 1.025 0.774 1.195 0.769 1.092 0.769 1.818 1.072 2.283 1.222
Avg 0.119 0.347 0.347 0.393 0.399 0.426 0.360 0.403 0.376 0.410 0.375 0.411 0.315 0.404 0.416 0.443 0.782 0.681 0.354 0.414 0.410 0.427 0.519 0.454 0.461 0.454 1.268 0.851 1.121 0.798
ILI 24 0.290 0.311 1.701 0.823 2.122 0.874 2.014 0.899 0.432 0.487 1.724 0.843 2.684 1.112 2.317 0.934 3.041 1.186 2.398 1.040 2.527 1.020 3.228 0.945 2.294 0.945 3.103 1.181 4.381 1.425
36 0.305 0.336 1.941 0.888 2.289 0.931 2.115 0.943 0.446 0.479 1.536 0.752 2.667 1.068 1.972 0.920 3.406 1.232 2.646 1.088 2.615 1.007 2.679 0.848 1.825 0.848 3.720 1.298 4.442 1.416
48 0.311 0.342 1.847 0.893 2.165 0.908 2.188 0.972 0.477 0.510 1.821 0.832 2.558 1.052 2.238 0.940 3.459 1.221 2.614 1.086 2.359 0.972 2.622 0.900 2.010 0.900 3.895 1.291 4.559 1.443
60 0.368 0.390 1.976 0.952 2.085 0.909 2.114 0.968 0.578 0.564 1.923 0.842 2.747 1.110 2.027 0.928 3.640 1.305 2.804 1.146 2.487 1.016 2.857 0.963 2.178 0.963 4.099 1.360 4.651 1.474
Avg 0.319 0.345 1.866 0.889 2.165 0.906 2.108 0.946 0.483 0.510 1.751 0.817 2.664 1.086 2.139 0.931 3.387 1.236 2.616 1.090 2.497 1.004 2.847 0.914 2.077 0.914 3.704 1.283 4.508 1.440
Count 58 4 0 0 1 1 1 0 0 0 0 0 0 0 0

### v2

@@ 4.4 Main results and analysis
ETTm1 96 0.142 0.263 0.264 0.330 0.320 0.357 0.334 0.368 0.351 0.370 0.320 0.351 0.319 0.358 0.324 0.361 0.316 0.362 0.338 0.375 0.428 0.444 0.345 0.372 0.375 0.398 0.379 0.419 0.386 0.398
192 0.162 0.282 0.316 0.368 0.361 0.381 0.377 0.391 0.392 0.393 0.375 0.376 0.369 0.384 0.362 0.383 0.363 0.390 0.374 0.387 0.445 0.468 0.380 0.389 0.408 0.410 0.426 0.441 0.459 0.444
336 0.174 0.295 0.315 0.378 0.390 0.404 0.426 0.420 0.424 0.413 0.411 0.399 0.402 0.406 0.390 0.402 0.408 0.426 0.410 0.411 0.533 0.519 0.413 0.413 0.435 0.428 0.445 0.459 0.495 0.464
720 0.184 0.304 0.382 0.425 0.454 0.441 0.491 0.459 0.485 0.448 0.464 0.431 0.464 0.442 0.461 0.438 0.481 0.476 0.478 0.450 0.728 0.655 0.474 0.453 0.499 0.462 0.543 0.490 0.585 0.516
Avg 0.166 0.286 0.319 0.375 0.381 0.395 0.407 0.410 0.413 0.406 0.393 0.389 0.389 0.398 0.384 0.396 0.392 0.414 0.400 0.406 0.534 0.522 0.403 0.407 0.429 0.425 0.448 0.452 0.481 0.456
ETTm2 96 0.111 0.219 0.183 0.265 0.175 0.258 0.180 0.264 0.181 0.264 0.181 0.264 0.175 0.258 0.177 0.260 0.179 0.275 0.187 0.267 0.197 0.321 0.193 0.292 0.189 0.280 0.203 0.287 0.192 0.274
192 0.129 0.236 0.248 0.305 0.237 0.299 0.250 0.309 0.246 0.304 0.242 0.302 0.240 0.230 0.248 0.306 0.307 0.376 0.249 0.309 0.326 0.375 0.284 0.362 0.253 0.319 0.269 0.328 0.280 0.339
336 0.140 0.247 0.311 0.345 0.298 0.340 0.311 0.348 0.306 0.341 0.302 0.341 0.301 0.338 0.304 0.342 0.325 0.388 0.321 0.351 0.372 0.421 0.369 0.427 0.314 0.357 0.325 0.366 0.334 0.361
720 0.152 0.255 0.101 0.396 0.391 0.396 0.412 0.407 0.407 0.397 0.403 0.396 0.404 0.398 0.403 0.397 0.502 0.490 0.408 0.403 0.410 0.448 0.554 0.522 0.414 0.413 0.421 0.415 0.417 0.413
Avg 0.133 0.239 0.211 0.328 0.275 0.323 0.288 0.332 0.285 0.327 0.282 0.326 0.280 0.306 0.283 0.326 0.328 0.382 0.291 0.333 0.326 0.391 0.350 0.401 0.293 0.342 0.305 0.349 0.306 0.347
ETTh1 96 0.203 0.326 0.338 0.368 0.375 0.400 0.386 0.405 0.381 0.391 0.369 0.387 0.381 0.391 0.394 0.408 0.398 0.427 0.384 0.402 0.429 0.440 0.386 0.400 0.494 0.479 0.376 0.419 0.513 0.491
192 0.228 0.347 0.377 0.385 0.429 0.421 0.441 0.436 0.434 0.422 0.414 0.419 0.433 0.420 0.446 0.438 0.430 0.453 0.436 0.429 0.494 0.482 0.437 0.432 0.538 0.504 0.420 0.448 0.534 0.504
336 0.238 0.355 0.400 0.404 0.484 0.458 0.487 0.458 0.474 0.446 0.451 0.438 0.472 0.441 0.485 0.455 0.460 0.460 0.491 0.469 0.706 0.625 0.481 0.459 0.574 0.521 0.459 0.465 0.588 0.535
720 0.245 0.363 0.465 0.439 0.498 0.482 0.503 0.491 0.464 0.463 0.483 0.475 0.485 0.471 0.495 0.474 0.491 0.509 0.521 0.500 0.750 0.689 0.519 0.516 0.562 0.535 0.506 0.507 0.643 0.616
Avg 0.229 0.348 0.395 0.399 0.447 0.440 0.454 0.448 0.438 0.431 0.429 0.430 0.443 0.431 0.455 0.444 0.445 0.462 0.458 0.450 0.595 0.559 0.456 0.452 0.542 0.510 0.440 0.460 0.570 0.537
ETTh2 96 0.147 0.259 0.385 0.396 0.289 0.341 0.297 0.349 0.290 0.339 0.292 0.341 0.289 0.338 0.302 0.348 0.332 0.377 0.340 0.374 0.632 0.547 0.333 0.387 0.340 0.391 0.358 0.397 0.476 0.458
192 0.168 0.279 0.438 0.425 0.372 0.392 0.380 0.400 0.375 0.388 0.377 0.392 0.375 0.391 0.388 0.400 0.422 0.441 0.402 0.414 0.876 0.663 0.477 0.476 0.430 0.439 0.429 0.439 0.512 0.493
336 0.192 0.295 0.451 0.437 0.386 0.414 0.428 0.432 0.414 0.425 0.419 0.429 0.425 0.435 0.426 0.433 0.447 0.474 0.452 0.452 0.924 0.702 0.594 0.541 0.485 0.479 0.496 0.487 0.552 0.551
720 0.223 0.310 0.470 0.461 0.412 0.434 0.427 0.445 0.419 0.437 0.424 0.441 0.435 0.449 0.431 0.446 0.442 0.467 0.462 0.468 1.390 0.863 0.831 0.657 0.500 0.497 0.463 0.474 0.562 0.560
Avg 0.183 0.286 0.436 0.430 0.364 0.395 0.383 0.407 0.375 0.397 0.378 0.401 0.381 0.403 0.387 0.407 0.411 0.440 0.414 0.427 0.956 0.694 0.559 0.515 0.439 0.452 0.437 0.449 0.526 0.516
Electricity 96 0.131 0.233 0.179 0.263 0.153 0.247 0.148 0.240 0.293 0.401 0.165 0.248 0.146 0.244 0.195 0.285 0.164 0.269 0.168 0.272 0.254 0.347 0.197 0.282 0.187 0.304 0.193 0.308 0.169 0.273
192 0.135 0.239 0.185 0.269 0.166 0.256 0.162 0.253 0.268 0.378 0.181 0.266 0.163 0.259 0.199 0.289 0.177 0.285 0.184 0.289 0.261 0.353 0.196 0.285 0.199 0.315 0.201 0.315 0.182 0.286
336 0.138 0.245 0.201 0.285 0.185 0.277 0.178 0.269 0.355 0.452 0.197 0.282 0.180 0.279 0.215 0.305 0.193 0.304 0.198 0.300 0.273 0.364 0.209 0.301 0.212 0.329 0.214 0.329 0.200 0.304
720 0.142 0.250 0.242 0.318 0.225 0.310 0.225 0.317 0.416 0.498 0.238 0.315 0.216 0.307 0.256 0.337 0.212 0.321 0.220 0.320 0.303 0.388 0.245 0.333 0.233 0.345 0.246 0.355 0.222 0.321
Avg 0.137 0.242 0.202 0.284 0.182 0.272 0.178 0.270 0.333 0.432 0.195 0.278 0.176 0.272 0.216 0.304 0.187 0.295 0.193 0.295 0.273 0.363 0.212 0.300 0.208 0.323 0.214 0.327 0.193 0.296
Traffic 96 0.228 0.286 0.376 0.265 0.462 0.285 0.395 0.268 0.898 0.572 0.463 0.297 0.483 0.321 0.544 0.359 0.519 0.309 0.593 0.321 0.558 0.320 0.650 0.396 0.607 0.392 0.587 0.366 0.612 0.338
192 0.239 0.292 0.392 0.271 0.473 0.296 0.417 0.276 0.763 0.522 0.469 0.298 0.490 0.321 0.540 0.354 0.537 0.315 0.617 0.336 0.572 0.331 0.598 0.370 0.621 0.399 0.604 0.373 0.613 0.340
336 0.243 0.300 0.405 0.277 0.498 0.296 0.433 0.283 0.894 0.608 0.484 0.305 0.506 0.330 0.551 0.358 0.534 0.313 0.629 0.336 0.587 0.342 0.605 0.373 0.622 0.396 0.621 0.383 0.618 0.328
720 0.256 0.307 0.436 0.294 0.506 0.313 0.467 0.302 1.019 0.646 - - 0.538 0.341 0.586 0.375 0.577 0.325 0.640 0.350 0.652 0.359 0.645 0.394 0.632 0.396 0.626 0.382 0.653 0.355
Avg 0.242 0.296 0.402 0.277 0.484 0.297 0.428 0.282 0.894 0.587 0.472 0.300 0.504 0.328 0.555 0.362 0.542 0.316 0.620 0.336 0.592 0.338 0.625 0.383 0.621 0.396 0.610 0.376 0.624 0.340
Exchange 96 0.097 0.209 0.083 0.196 0.100 0.222 0.086 0.206 0.089 0.210 0.083 0.201 0.082 0.199 0.082 0.201 0.102 0.235 0.107 0.234 0.329 0.440 0.088 0.218 0.085 0.204 0.148 0.237 0.111 0.237
192 0.103 0.217 0.170 0.291 0.212 0.326 0.177 0.299 0.182 0.303 0.175 0.296 0.176 0.297 0.187 0.307 0.172 0.316 0.226 0.344 0.544 0.586 0.176 0.315 0.182 0.303 0.271 0.335 0.219 0.335
336 0.117 0.232 0.316 0.404 0.379 0.442 0.331 0.417 0.327 0.414 0.327 0.413 0.337 0.418 0.345 0.427 0.272 0.407 0.367 0.448 1.017 0.786 0.313 0.427 0.348 0.428 0.460 0.476 0.421 0.476
720 0.158 0.279 0.817 0.679 0.904 0.715 0.847 0.691 0.852 0.696 0.849 0.694 0.909 0.725 0.887 0.708 0.714 0.658 0.964 0.746 1.239 0.912 0.839 0.695 1.025 0.774 1.195 0.769 1.092 0.769
Avg 0.119 0.234 0.347 0.393 0.399 0.426 0.360 0.403 0.363 0.406 0.359 0.401 0.376 0.410 0.375 0.411 0.315 0.404 0.416 0.443 0.782 0.681 0.354 0.414 0.410 0.427 0.519 0.454 0.461 0.454
ILI 24 0.290 0.311 1.701 0.823 2.122 0.874 2.014 0.899 3.340 1.299 2.585 1.065 0.432 0.487 1.724 0.843 2.684 1.112 2.317 0.934 3.041 1.186 2.398 1.040 2.527 1.020 3.228 0.945 2.294 0.945
36 0.305 0.336 1.941 0.888 2.289 0.931 2.115 0.943 4.016 1.453 2.623 1.085 0.446 0.479 1.536 0.752 2.667 1.068 1.972 0.920 3.406 1.232 2.646 1.088 2.615 1.007 2.679 0.848 1.825 0.848
48 0.311 0.342 1.847 0.893 2.165 0.908 2.188 0.972 4.541 1.554 2.465 1.035 0.477 0.510 1.821 0.832 2.558 1.052 2.238 0.940 3.459 1.221 2.614 1.086 2.359 0.972 2.622 0.900 2.010 0.900
60 0.368 0.390 1.976 0.952 2.085 0.909 2.114 0.968 4.406 1.517 2.386 1.029 0.578 0.564 1.923 0.842 2.747 1.110 2.027 0.928 3.640 1.305 2.804 1.146 2.487 1.016 2.857 0.963 2.178 0.963
Avg 0.319 0.345 1.866 0.889 2.165 0.906 2.108 0.946 4.076 1.456 2.515 1.054 0.483 0.510 1.751 0.817 2.664 1.086 2.139 0.931 3.387 1.236 2.616 1.090 2.497 1.004 2.847 0.914 2.077 0.914
Count 58 5 0 0 0 0 1 1 0 0 0 0 0 0 0
Short-term forecasting. Short-term forecasting results for the time series are presented in Table 3. Analysis of this table yields the following conclusions:
  • The predictions generated by ACNet surpass those of other baseline models, with the Transformer-based Crossformer model achieving the second-best results. The success of ACNet can be attributed to its comprehensive feature extraction approach, effectively capturing and utilizing local, global, and nonlinear features. This significantly enhances the model’s predictive performance. Crossformer also demonstrates its effectiveness in capturing complex patterns within real-world time series by explicitly extracting the interdependencies between different variables in the time series.
  • Transformer-based models (e.g., iTransformer, PatchTST) exhibit limited capabilities in handling time series forecasting tasks involving nonlinear features. This limitation may stem from their reliance on the self-attention mechanism to capture global information from the input sequence. Nonlinear relationships, however, often require local contextual information for better understanding and accurate feature extraction.
  • The MICN model, which is based on multi-scale dilated convolution, shows the poorest performance across all datasets. This outcome suggests that traditional convolutions with fixed-size receptive fields are inadequate for real-world time series datasets containing numerous nonlinear features. The fixed nature of these receptive fields restricts their flexibility in handling nonlinear features of varying scales, leading to information redundancy and diminished predictive performance.
Table 3: Short-term multivariate forecasting results with different metrics. The best results are in bold numbers and the second best are highlighted with an underline.
Models
ACNet
ConvTimeNet
TimeMixer
iTransformer
FITS
PDF
TSLANet
TSMixer
PatchTST
MICN
TimesNet
Crossformer
PEMS03 MAE 0.145 0.245 0.187 0.178 0.348 0.195 0.185 0.198 0.212 0.462 0.190 0.166
MAPE 1.295 1.898 1.466 1.364 2.656 2.794 1.459 1.567 1.636 2.948 1.455 1.321
RMSE 0.208 0.361 0.277 0.266 0.468 0.293 0.278 0.277 0.310 0.606 0.290 0.253
PEMS04 MAE 0.183 0.307 0.247 0.235 0.448 0.260 0.272 0.238 0.269 0.935 0.269 0.206
MAPE 1.550 2.635 2.039 1.931 3.658 2.131 2.179 1.985 2.177 7.799 2.208 1.682
RMSE 0.311 0.566 0.487 0.481 0.623 0.509 0.526 0.434 0.516 1.298 0.534 0.430
PEMS07 MAE 0.130 0.257 0.187 0.181 0.382 0.194 0.197 0.186 0.211 0.622 0.199 0.150
MAPE 1.349 2.561 1.907 1.849 3.509 2.004 1.948 1.970 2.058 4.773 1.904 1.667
RMSE 0.191 0.368 0.279 0.274 0.501 0.290 0.297 0.273 0.303 0.800 0.302 0.232
PEMS08 MAE 0.124 0.313 0.259 0.255 0.473 0.268 0.256 0.259 0.285 0.797 0.302 0.238
MAPE 1.594 2.115 1.649 1.705 3.235 1.704 1.709 1.685 1.800 5.239 2.077 1.497
RMSE 0.204 0.582 0.523 0.516 0.723 0.531 0.517 0.481 0.540 1.189 0.664 0.497